A metal of length 15.01m is heated until its temperature rises by 600C. If its new length is 15.05, calculate its linear expansivity


1
Expert's answer
2021-03-24T19:38:49-0400

The linear expansivity is


"\alpha=\frac{L_2-L_1}{L_1\Delta T}=\frac{15.05-15.01}{15.01·600}=4.44·10^{-6}\text{ K}^{-1}."